# Gökdeniz Gülmez Machine Learning Research Stuttgart, Germany

### v2

@@ Abstract
We present Gabliteration, a novel neural weight modification technique that advances beyond traditional abliteration methods by implementing adaptive multi-directional projections with regularized layer selection. Our approach addresses the fundamental limitation of existing methods that compromise model quality while attempting to modify specific behavioral patterns. Through dynamic layer optimization, regularized projection matrices, and adaptive scaling mechanisms, we achieve theoretically superior weight modification while minimizing quality degradation in unrelated domains. We validate our method through the gabliterated-v1 model series (0.6B to 32B parameters) available on Hugging Face [gulmezhf], demonstrating practical applicability across multiple model scales.
Keywords: Large Language Models, Neural Network Modification, Weight Ablation, Behavioral Alignment, Transformer Architecture
@@ 1 Introduction
The field of neural network behavioral modification has witnessed significant developments in recent years, particularly following the groundbreaking work by Arditi et al. (2024) [arditi2024refusal] on refusal direction identification in large language models. Their paper ‘Refusal in Language Models Is Mediated by a Single Direction’ established that behavioral patterns in language models can be effectively modified through targeted weight ablation techniques, which they termed "abliteration".
The practical implementation of these concepts followed a rapid community-driven evolution. After Arditi et al [arditi2024refusal]. first published their findings, Maxime Labonne (mlabonne [mlabonne2025models]) validated the concept with a real-world implementation using the Llama3 model. This was followed by our viral ‘Josiefied-abliterated-Qwen2.5’ [gulmez2024qwen25] implementation, which demonstrated the technique’s potential for widespread adoption. Subsequently, huihui.ai [huihuiai2025models] applied these methods to their models using huggingface’s transformers library, after which we further advanced the approach with Qwen3 [gulmez2025qwen3].
@@ 2 Mathematical Foundation
Building upon the foundation established by Arditi et al. [arditi2024refusal], we extend their single-direction approach to a comprehensive multi-directional framework with theoretical guarantees.
@@ 2.1 Multi-Directional Refusal Vector Extraction
While Arditi et al. [arditi2024refusal] demonstrated the effectiveness of single-direction refusal vectors, we hypothesized that behavioral patterns exist in higher-dimensional subspaces. Our multi-directional extraction employs singular value decomposition (SVD) on a paired difference matrix constructed between harmful and harmless representations.
@@ 2.7 Relationship to Prior Orthogonalization Methods
The original abliteration [arditi2024refusal] uses a rank-1 projection:
@@ 3.4 Implementation Notes
    Batch Processing: For efficiency, hidden state extraction should process prompts in batches of size $b\approx 8$-1616 depending on GPU memory.
@@
    Refusal Keywords: The set $\mathcal{R}$ should include common refusal patterns: {"I cannot", "I apologize", "I’m unable", "I can’t assist", "I cannot provide", "I’m not able", "I cannot help"}.
@@ 4.2 Dataset Construction
We successfully applied the Gabliteration technique to create the base Gabliterated-v1 and JOSIEfied-Gabliterated-v1 model series, available on our Hugging Face profile (Goekdeniz-Guelmez). These models demonstrate consistent improvements across multiple scales, validating the scalability and robustness of the proposed approach. They achieve targeted behavioral modification while maintaining strong performance preservation across standard benchmarks, as documented on their respective model pages on Hugging Face [gulmezhf].
@@ 5.1 Performance Preservation
Assumptions: The bound assumes that weight matrices admit a decomposition $\mathbf{W}^{(\ell)}=\mathbf{W}_{\mathcal{T}}^{(\ell)}+\mathbf{W}_{\mathcal{R}}^{(\ell)}+\mathbf{W}_{\perp}^{(\ell)}$, where components are approximately orthogonal. A formal proof is provided in Appendix .3.
@@ 7.3 Ablation Study: Pairing Methods
Results (Table 1): Across all eight evaluated models, Gabliteration (Ours) exhibited lower average KL divergence, indicating reduced distributional drift and better preservation of the original model behavior while lovering the average refusal rates as well. Abliteration showed significantly higher variance across models, suggesting reduced robustness and sensitivity to model scale and architecture.
Table 1: Per-model refusal rate and KL divergence for Gabliteration and Abliteration. All models are evaluated on identical datasets (mlabonne/harmful_behaviors vs. mlabonne/harmless_alpaca), using the same refusal keywords, 100 trials, 100 harmful evaluation prompts, and 400 samples for direction extraction. Abliteration has been applied using the heretic package by Philipp Emanuel Weidmann [heretic]. Lower is better for both metrics.
@@
Qwen/Qwen3–4B-Instruct-2507 4 0.1522 21 0.4300
Qwen/Qwen3–4B-Thinking-2507 3 0.0014 4 0.0500
@@
Nanbeige/Nanbeige4–3B-Thinking-2511 8 0.0074 11 0.0024
@@ Acknowledgments
We acknowledge the foundational work of Arditi et al. [arditi2024refusal] on single-direction abliteration, which provided the initial inspiration for developing our Gabliteration technique. We are grateful to Joshua Ollswang, Awni Hannun for their thoughtful feedback, which contributed to the improvement of this work.
.2 Weight Modification Bound Proof
Theorem 1: Under the assumption that the refusal directions $\mathbf{r}_{i}$ span a low-dimensional subspace and that the regularization parameter $\lambda>0$, the single-pass Gabliteration modification in Phase 5 satisfies:
$\|\mathbf{W}_{\text{modified}}-\mathbf{W}_{\text{original}}\|_{F}\leq\sum_{\ell\in\mathcal{L}_{\text{eff}}}\alpha_{\ell}\|\mathbf{W}_{\ell}\mathbf{P}\|_{F}$
where $\alpha_{\ell}$ is the adaptive scaling factor at layer $\ell$, $\mathbf{P}$ is the regularized projection matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
Proof: At each modified layer $\ell\in\mathcal{L}_{\text{eff}}$, the weight update is applied exactly once:
$\mathbf{W}_{\text{modified}}^{(\ell)}=\mathbf{W}_{\text{original}}^{(\ell)}-\alpha_{\ell}\left(\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\right)$
The modification magnitude at layer $\ell$ is:
$\|\mathbf{W}_{\text{modified}}^{(\ell)}-\mathbf{W}_{\text{original}}^{(\ell)}\|_{F}=\|\alpha_{\ell}\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}=\alpha_{\ell}\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}$
Since modifications are applied independently to each layer (Phase 5 iterates over layers but applies each modification once), and unmodified layers contribute zero perturbation, the squared Frobenius norm of the total modification satisfies:
$\|\mathbf{W}_{\text{modified}}-\mathbf{W}_{\text{original}}\|_{F}^{2}=\sum_{\ell\in\mathcal{L}_{\text{eff}}}\|\mathbf{W}_{\text{modified}}^{(\ell)}-\mathbf{W}_{\text{original}}^{(\ell)}\|_{F}^{2}$
This follows from the fact that modifications at different layers are independent and affect disjoint weight matrices. Taking the square root and applying the Cauchy-Schwarz inequality:
$\|\mathbf{W}_{\text{modified}}-\mathbf{W}_{\text{original}}\|_{F}=\sqrt{\sum_{\ell\in\mathcal{L}_{\text{eff}}}\alpha_{\ell}^{2}\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}^{2}}$
By the triangle inequality, specifically noting that $\sqrt{\sum_{i}a_{i}^{2}}\leq\sum_{i}|a_{i}|$ for non-negative terms $a_{i}\geq 0$:
$\|\mathbf{W}_{\text{modified}}-\mathbf{W}_{\text{original}}\|_{F}=\sqrt{\sum_{\ell\in\mathcal{L}_{\text{eff}}}\alpha_{\ell}^{2}\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}^{2}}\leq\sum_{\ell\in\mathcal{L}_{\text{eff}}}\alpha_{\ell}\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}$
[…]
Key Assumptions for Theorem 1: 1. Weight decomposition: $\mathbf{W}^{(\ell)}=\mathbf{W}_{\mathcal{T}}^{(\ell)}+\mathbf{W}_{\mathcal{R}}^{(\ell)}+\mathbf{W}_{\perp}^{(\ell)}$ 2. Projection bound: $\|\mathbf{W}_{\mathcal{T}}\mathbf{P}\|_{F}\leq\frac{\|\mathbf{R}\|_{F}^{2}}{\|\mathbf{R}\|_{F}^{2}+\lambda}\|\mathbf{W}_{\mathcal{T}}\|_{F}\cos(\theta)$ 3. Small regularization: $\lambda\ll\|\mathbf{R}\|_{F}^{2}$
Remark: This theorem characterizes the single-pass nature of Algorithm 1, where weights are modified exactly once per layer during Phase 5. The bound depends on:
  • The adaptive scaling factors $\alpha_{\ell}$, which vary by layer position (Section 2.5). Middle layers receive higher scaling ($\alpha_{\text{middle}}\approx\alpha_{\text{base}}(1+\beta)$), while boundary layers receive lower scaling ($\alpha_{\text{boundary}}\approx\alpha_{\text{base}}$).
  • The projection strength $\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}$, which measures how much of each weight matrix aligns with the refusal subspace $\mathcal{R}$.
  • The number of effective layers $|\mathcal{L}_{\text{eff}}|$, determined by the threshold $\tau$ in Phase 4. Empirically, $|\mathcal{L}_{\text{eff}}|\approx 0.23L$ for typical choices of $\tau=0.8$.
Corollary (Bounded Total Modification): If $\|\mathbf{W}_{\text{original}}^{(\ell)}\mathbf{P}\|_{F}\leq C$ for all $\ell$ and $\alpha_{\ell}\leq\alpha_{\max}=\alpha_{\text{base}}(1+\beta)$, then:
$\|\mathbf{W}_{\text{modified}}-\mathbf{W}_{\text{original}}\|_{F}\leq|\mathcal{L}_{\text{eff}}|\cdot\alpha_{\max}\cdot C$
This corollary provides a concrete upper bound in terms of hyperparameters, showing that the total modification scales linearly with the number of effective layers and the maximum scaling factor.
.3 Performance Preservation Bounds

### v3

@@ Abstract
We introduce Gabliteration, a principled and mathematically rigorous neural weight modification framework that substantially advances beyond conventional abliteration paradigms through the systematic incorporation of adaptive multi-directional projections with theoretically motivated layer-wise regularization. Our methodology addresses the critical and previously unresolved limitation endemic to existing behavioral modification techniques. The inherent trade-off between selective behavioral alteration and the preservation of general model fidelity across unrelated downstream tasks. By leveraging dynamic layer selection mechanisms informed by distribution-aware separability metrics, coupled with multi-directional singular value decomposition-based direction extraction and adaptive scaling through regularized projection matrices, we achieve theoretically justified weight modification strategies that demonstrably attenuate performance degradation in semantically orthogonal domains. Furthermore, we provide rigorous theoretical analysis establishing convergence guarantees and approximation bounds for our proposed optimization formulation. Our empirical validation spans the gabliterated model series, comprehensively available on Hugging Face [5], thereby establishing both practical scalability and reproducibility across diverse model architectures and parameter regimes. The results demonstrate statistically significant improvements in behavioral selectivity while preserving generative quality, positioning Gabliteration as a robust and generalizable framework for controlled behavioral alignment in large-scale language models.
Keywords: Large Language Models, Neural Network Modification, Targeted Weight Ablation, Behavioral Alignment
@@ 1 Introduction
The field of neural network behavioral modification has witnessed significant developments in recent years, particularly following the groundbreaking work by Arditi et al. (2024) [1] on refusal direction identification in large language models. Their paper ‘Refusal in Language Models Is Mediated by a Single Direction’ established that behavioral patterns in language models can be effectively modified through targeted weight ablation techniques, which they termed ‘abliteration’.
The practical implementation of these concepts followed a rapid community-driven evolution. After Arditi et al [1]. first published their findings, Maxime Labonne (mlabonne [14]) validated the concept with a real-world implementation using the Llama3 model. This was followed by our viral ‘Josiefied-abliterated-Qwen2.5’ [4] implementation, which demonstrated the technique’s potential for widespread adoption. Subsequently, huihui.ai [13] applied these methods to their models using huggingface’s transformers library, after which we further advanced the approach with Qwen3 [6].
@@ 2 Mathematical Foundation
Building upon the foundation established by Arditi et al. [1], we extend their single-direction approach to a comprehensive multi-directional framework with theoretical guarantees.
@@ 2.1 Multi-Directional Refusal Vector Extraction
While Arditi et al. [1] demonstrated the effectiveness of single-direction refusal vectors, we hypothesized that behavioral patterns exist in higher-dimensional subspaces. Our multi-directional extraction employs singular value decomposition (SVD) on a paired difference matrix constructed between harmful and harmless representations.
@@
2.2 Theoretical Foundation: Neural Superposition and Multi-Directional Extraction
The multi-directional extraction framework employed by Gabliteration is grounded in recent theoretical advances in understanding how neural networks represent features in their hidden states. The work by Elhage et al. (2022) [3] on toy models of superposition has demonstrated a fundamental principle: neural networks do not represent individual features in isolated dimensions, but rather can encode multiple distinct features within the same representational subspace through a phenomenon known as superposition. This insight has profound implications for understanding behavioral modification in large language models.
2.2.1 Feature Superposition in Neural Networks
In dense, high-dimensional neural network representations, many distinct behavioral patterns can coexist in non-orthogonal subspaces, overlapping in the same hidden dimensions. A single behavioral direction (such as those identified by Arditi et al.) may capture only the primary axis of variation, leaving substantial behavioral structure unrepresented in orthogonal dimensions. This is precisely the regime in which superposition becomes prevalent: when features (behavioral patterns) are sparse and the network can benefit from representing more features than it has explicit dedicated dimensions.
When we extract multiple directions via singular value decomposition on the harmful-harmless difference matrix, we are fundamentally discovering the low-rank structure of the behavioral subspace. The top singular vectors correspond to the principal axes along which harmful and harmless representations diverge. By extracting $k>1$ directions, we capture not only the primary refusal axis but also secondary, tertiary, and higher-order behavioral signatures that exist in superposition with the primary direction.
2.2.2 Implications for Behavioral Modification
The superposition principle explains why modification along a single direction can inadvertently disrupt other, seemingly unrelated capabilities. When behavioral features exist in superposition, a naive single-direction projection removes not just the target behavior but also disrupts the shared geometric substrate that supports multiple overlapping functions. These overlapping functions may include legitimate model capabilities needed for general knowledge retrieval, reasoning, and generation.
Gabliteration’s multi-directional approach specifically addresses this problem by:
  1. 1.
    Identifying the full behavioral subspace: Rather than assuming refusal behavior lies along a single dimension, we extract the dominant $k$ directions of behavioral divergence. The lower singular values indicate which directions contain genuine refusal signal versus noise, allowing us to distinguish robust behavioral patterns from artifacts.
  2. 2.
    Minimal footprint modification: By orthogonally projecting weight matrices onto only the identified behavioral subspace, we ensure that modifications affect the narrowest possible region of the neural manifold. This minimizes collateral damage to unrelated features that may reside in nearby but distinct geometric regions.
  3. 3.
    Preserving feature geometry: The regularized projection mechanism (detailed in Section 2.3) maintains the geometric structure of the weight matrices while removing only the projections aligned with behavioral directions. This approach respects the underlying feature geometry learned during pretraining, preventing the model from having to re-learn unrelated capabilities.
  4. 4.
    Exploiting sparsity of behavioral directionality: Empirically, harmful and harmless behavioral divergence is sparse in the sense that it concentrates in a small number of dominant directions (typically 1–3 singular vectors explain $\geq 90\%$ of the divergence). This sparsity implies that the refusal subspace has low intrinsic dimensionality, allowing accurate recovery with $k$ much smaller than the full hidden dimension $d$. This is consistent with the sparsity principle that underlies superposition in dense representations.
2.2.3 Connection to Compressed Sensing
The superposition framework also connects to principles from compressed sensing: sparse signals in high-dimensional spaces can be recovered from low-dimensional measurements without loss of information. In our setting, the behavioral signal is sparse (concentrated in a small subspace), the neural representation is high-dimensional ($d\sim 4096$ for typical modern LLMs), and the measurement process (hidden state computation) is efficient. Consequently, extracting multiple behavioral directions via SVD achieves stable recovery of the behavioral subspace with empirical sample complexity that scales gracefully with $k$ rather than with the full dimension $d$.
2.2.4 Why Multi-Directional Extraction Matters
The standard abliteration approach (single-direction) implicitly assumes that behavioral modification is one-dimensional—a strong assumption that the superposition perspective shows to be often false. Empirically (as demonstrate in Elhage et al. [3]), harmful and harmless behaviors diverge along multiple distinct but correlated directions. A single-direction intervention leaves residual refusal capacity along secondary directions, reducing modification effectiveness. Conversely, an over-aggressive modification along a single direction may over-correct in directions where the target behavior is less pronounced, causing unnecessary collateral damage.
The multi-directional formulation elegantly balances these concerns: it discovers the intrinsic dimensionality of behavioral divergence through SVD and targets only the statistically significant directions (as ranked by singular value magnitude). This data-driven approach inherits the geometric interpretability of single-direction methods while recovering the flexibility of higher-dimensional subspace approaches.
@@ 2.5 Dynamic Layer Selection Algorithm
While effective as a first-order signal, this criterion has several fundamental limitations that should be acknowledged in the next version of Gabliteration:
First, the metric is scale-sensitive. Hidden state norms vary substantially across depth in transformer models, which biases selection toward layers with large activation magnitudes rather than layers with genuinely informative behavioral separation. As a result, layers with high energy but weak discriminative structure may be preferred over layers with cleaner but lower-norm separation.
Second, the metric captures only translational separation and ignores angular structure. Two layers may exhibit similar mean differences in norm while encoding harmful and harmless behaviors in nearly parallel directions, making projection-based interventions ineffective or destructive. Since weight projection operates on directional geometry, ignoring angular dissimilarity leads to suboptimal layer selection.
Third, the approach does not account for signal-to-noise characteristics. A large mean difference can arise from high variance or unstable representations rather than a robust, low-noise refusal signal, increasing sensitivity to sampling noise and dataset composition.
Finally, the method provides no indication of cross-layer directional stability. It treats layers independently and cannot distinguish between isolated local effects and refusal directions that persist coherently across multiple consecutive layers. This limits the ability to safely apply extracted directions to neighboring or preceding layers.
@@ 2.8 Relationship to Prior Orthogonalization Methods
The original abliteration [1] uses a rank-1 projection:
@@ 3.4 Implementation Notes
    Batch Processing: For efficiency, hidden state extraction should process prompts in batches of size $b\approx 8$-$16$ depending on GPU memory.
@@
    Refusal Keywords: The set $\mathcal{R}$ should include common refusal patterns: {"I cannot", "I apologize", "I’m unable", …}.
@@ 4.2 Dataset Construction
We successfully applied the Gabliteration technique to create the base Gabliterated-v1 and JOSIEfied-Gabliterated-v1 model series, available on our Hugging Face profile (Goekdeniz-Guelmez). These models demonstrate consistent improvements across multiple scales, validating the scalability and robustness of the proposed approach. They achieve targeted behavioral modification while maintaining strong performance preservation across standard benchmarks, as documented on their respective model pages on Hugging Face [5].
@@ 5.1 Performance Preservation
Assumptions: The bound assumes that weight matrices admit a decomposition $\mathbf{W}^{(\ell)}=\mathbf{W}_{\mathcal{T}}^{(\ell)}+\mathbf{W}_{\mathcal{R}}^{(\ell)}+\mathbf{W}_{\perp}^{(\ell)}$, where components are approximately orthogonal. A formal proof is provided in Appendix A.2.
@@ 7.3 Ablation Study: Pairing Methods
Results (Table 1): Across all ten evaluated models, Gabliteration (Ours) exhibited lower average KL divergence, indicating reduced distributional drift and better preservation of the original model behavior while lovering the average refusal rates as well. Abliteration showed significantly higher variance across models, suggesting reduced robustness and sensitivity to model scale and architecture.
Table 1: Per-model refusal rate and KL divergence for Gabliteration and Abliteration. All models are evaluated on identical datasets (mlabonne/harmful_behaviors vs. mlabonne/harmless_alpaca), using the same refusal keywords, 100 trials, 100 harmful evaluation prompts, and 400 samples for direction extraction. Abliteration has been applied using the heretic package by Philipp Emanuel Weidmann [16]. Lower is better for both metrics.
@@
Qwen/Qwen3–4B-Instruct–2507 4 0.2522 21 0.4300
Qwen/Qwen3–4B-Thinking–2507 2 0.1140 4 0.0500
@@
Nanbeige/Nanbeige4–3B-Thinking–2511 8 0.1110 11 0.0024
allenai/Olmo–3–7B-Instruct 17 0.0018 31 0.1400
ZeroXClem/Qwen3–4B-Sky-High-Hermes 2 0.0992 6 0.3691
@@ 8.2 Future Work
9 Reproducibility and Package Availability
To facilitate reproducibility and enable researchers and practitioners to apply Gabliteration independently, we have developed and released the Gabliteration Python package [7]. The package is publicly available on GitHub at https://github.com/Goekdeniz-Guelmez/gabliteration and provides a complete, production-ready, fully automated command-line interface (CLI) for applying the Gabliteration algorithm to any HuggingFace-compatible language model.
The package includes:
  • Fully automated CLI with no manual configuration required for basic usage.
  • Complete implementation of the Gabliteration algorithm with all components (multi-directional refusal vector extraction, ridge-regularized projection, dynamic layer selection, and adaptive scaling).
  • Automatic hidden state extraction, separability metric computation, and layer effectiveness evaluation.
  • Seamless support for multible transformer architectures available on HuggingFace (Qwen, Llama, Mistral, and others).
  • Configurable hyperparameters (number of gabliterated versions, test samples, batch size, KL-divergence samples, max tokens, etc.) with optimized defaults.
  • Automated model saving and output management.
Users can now apply Gabliteration to any model with a single command line invocation, making the methodology immediately accessible to the broader research community. This enables independent verification of our results and facilitates widespread adoption of the technique. We encourage researchers to use this package, apply it to their own models, contribute improvements, and share their findings with the community.
@@ Acknowledgments
We acknowledge the foundational work of Arditi et al. [1] on single-direction abliteration, which provided the initial inspiration for developing our Gabliteration technique. We are grateful to Joshua Ollswang, Awni Hannun, and DontPlanToEnd for their thoughtful feedback, help, and evaluations, which contributed to the improvement of this work.
Appendix A Benchmark Results
A.1 UGI Benchmark Evaluation
We evaluated the effectiveness of Gabliteration on five Models:
  • Goekdeniz-Guelmez/Qwen3–4B-Sky-High-Hermes-gabliterated [11]
  • Goekdeniz-Guelmez/Qwen3–4B-Thinking-2507-gabliterated [12]
  • Goekdeniz-Guelmez/Qwen3–4B-Instruct-2507-gabliterated [10]
  • Goekdeniz-Guelmez/Gemma-3–1b-it-gabliterated [8]
  • Goekdeniz-Guelmez/Nanbeige4–3B-Thinking-2511-gabliterated [9]
using the UGI (Uncensored General Intelligence) benchmark suite [15]. The UGI benchmark evaluates models across multiple dimensions, including overall intelligence (UGI score), writing quality, natural intelligence (NatInt), and political bias. Notably, our gabliterated Qwen3 4B Thinking 2507 and Qwen3 4B Sky High Hermes achieved exceptional performance. Qwen3 4B Thinking 2507 ranked 10th worldwide, while Qwen3 4B Sky High Hermes reached an impressive 4th place in the W/10 score, making them the first 4B-parameter models to place within the global top ten, competing directly with models in the 32B-parameter and upper class [2].
A.1.1 Qwen3–4B-Sky-High-Hermes-Gabliterated
Table 2 reports the UGI benchmark results for the gabliterated Qwen3–4B-Sky-High-Hermes model. Despite the absence of a corresponding base-model benchmark, the gabliterated variant achieves the strongest UGI score of 33.95 and a W/10 score of 9.8 observed in any four billion parameter model, placing it competitively alongside substantially larger models. Writing quality (25.44) and natural intelligence (11.86) remain at high levels, while the political lean (-21.0%) indicates a stable ideological profile.
Table 2: UGI benchmarks: Qwen3–4B-Sky-High-Hermes-Gabliterated. Higher scores are better for UGI, W/10, Writing, and NatInt. Political lean indicates ideological positioning (negative values indicate left-leaning, and positive values indicate right-leaning respectively).
Metric Base Model Gabliterated Model
UGI Score 33.95
W/10 Score 9.8
Writing 25.44
NatInt 11.86
Political Lean (%) -21.0
Figure 1: UGI Leaderboard Ranking. The gabliterated Qwen3–4B-Sky-High-Hermes model ranks among substantially larger models, achieving a W/10 score of 9.8 despite its 4B parameter budget. This represents the first ever 4B-parameter model to achieve such performance on the UGI benchmark.
These results suggest that gabliteration can produce high-performing models even when applied to style-optimized variants with explicit reasoning specialization. The model is publicly available at Goekdeniz-Guelmez/Qwen3--4B-Sky-High-Hermes-gabliterated on Hugging Face [11].
A.1.2 Qwen3–4B-Thinking-2507-Gabliterated
Table 3 presents a detailed comparison between the base Qwen3–4B-Thinking–2507 model and its gabliterated variant. The gabliterated model demonstrates substantial improvements across key metrics, with a 55% increase in UGI score (from 20.78 to 32.25) and a 239% improvement in W/10 score (from 2.8 to 9.5), while maintaining comparable performance on writing quality and natural intelligence metrics. Notably, the political lean remains nearly unchanged (-26.0% vs -26.1%), indicating that Gabliteration successfully preserves the model’s ideological characteristics while enhancing its overall capabilities.
Table 3: UGI Benchmark Comparison: Qwen3–4B-Thinking–2507 Base vs Gabliterated.
Metric Base Model Gabliterated Model
UGI Score 20.78 32.25
W/10 Score 2.8 9.5
Writing 10.87 11.3
NatInt 16.16 16.67
Political Lean (%) -26.1 -26.0
Figure 2: UGI Leaderboard Ranking. The gabliterated Qwen3–4B-Thinking–2507 model ranks in the top 10 globally, achieving a W/10 score of 9.5 alongside models in the 32B parameter class as well.
The exceptional performance of the gabliterated Qwen3–4B-Thinking–2507 model demonstrates the practical effectiveness of our approach in enhancing model capabilities while maintaining parameter efficiency. The model is publicly available at Goekdeniz-Guelmez/Qwen3--4B-Thinking-2507-gabliterated [12] on Hugging Face. As shown in Figure 2, our model’s placement in the UGI leaderboard validates its exceptional performance across the benchmark suite.
A.1.3 Qwen3–4B-Instruct–2507–Gabliterated
Table 4 compares the base Qwen3–4B-Instruct–2507 model with its gabliterated variant. The gabliterated model exhibits a substantial improvement in UGI score, increasing by 113% from 13.1 to 27.92. The W/10 score improves by 143% (from 3.5 to 8.5), indicating a marked enhancement in reasoning-related performance.
Table 4: UGI Benchmark Comparison: Qwen3–4B-Instruct–2507 Base vs Gabliterated.
Metric Base Model Gabliterated Model
UGI Score 13.1 27.92
W/10 Score 3.5 8.5
Writing 29.92 27.03
NatInt 13.76 13.9
Political Lean (%) -12.4 -17.7
While writing quality shows a moderate decrease (from 29.92 to 27.03), natural intelligence remains effectively unchanged (13.76 to 13.9). The political lean shifts moderately from -12.4% to -17.7%, remaining within the same ideological region. These results demonstrate that gabliteration generalizes effectively to instruction-tuned models, yielding large gains in core benchmark performance without destabilizing overall model behavior. The model is available at Goekdeniz-Guelmez/Qwen3--4B-Instruct--2507--gabliterated [10].
A.1.4 Gemma–3–1B-IT-Gabliterated
Table 5 presents the UGI benchmark comparison for Gemma–3–1B-IT and its gabliterated variant. Despite a decrease in overall UGI score (from 15.03 to 13.49, -10.2%), the gabliterated model demonstrates a substantial improvement in W/10 score, increasing by 83% from 3.0 to 5.5. Writing quality also improves by 12.6%, rising from 24.01 to 27.03.
The political lean undergoes a significant shift from -19.3% to near-neutral (0.1%), indicating that gabliteration can meaningfully alter ideological positioning in smaller models. While NatInt could not be reliably measured for the gabliterated variant, the observed gains in reasoning efficiency and writing quality suggest that gabliteration remains effective even in the 1B parameter regime, albeit with stronger trade-offs than observed in larger models. The model is publicly available at Goekdeniz-Guelmez/Gemma--3--1B-IT-gabliterated [8].
Table 5: UGI Benchmark Comparison: Gemma–3–1B-IT Base vs Gabliterated.
Metric Base Model Gabliterated Model
UGI Score 15.03 13.49
W/10 Score 3 5.5
Writing 24.01 27.03
NatInt 11.19
Political Lean (%) -19.3 0.1
These results highlight the scalability of gabliteration across model sizes, showing that even lightweight instruction-tuned models can achieve near state-of-the-art reasoning efficiency without sacrificing stylistic or ideological stability. The model is publicly available at Goekdeniz-Guelmez/Gemma--3--1B-IT-gabliterated [8].
A.1.5 Nanbeige4–3B-Thinking–2511-Gabliterated
We further validated the generalization of Gabliteration on a smaller model scale by evaluating the Nanbeige4–3B-Thinking–2511 base model. This evaluation demonstrates that our approach maintains effectiveness across diverse model architectures and parameter scales. The gabliterated variant achieved notable improvements in UGI and W/10 scores while showing substantial enhancements in writing quality compared to the base model.
Table 6 presents the comparative results for the Nanbeige4–3B-Thinking–2511 model. The gabliterated version shows a 23% improvement in UGI score (from 21.16 to 26.05) and a 33% improvement in W/10 score (from 4.5 to 6.0), with a particularly notable 38% increase in writing quality (from 9.76 to 13.48). The natural intelligence metric also improves slightly (from 13.72 to 16.82), while the political lean shifts from -16.9% to -10.2%, indicating a moderate rightward shift in the model’s ideological positioning post-gabliteration.
Table 6: UGI Benchmark Comparison: Nanbeige4–3B-Thinking–2511 Base vs Gabliterated. The gabliterated model demonstrates consistent improvements in UGI, W/10, and Writing metrics, validating Gabliteration’s effectiveness across different model architectures. Higher scores are better for UGI, W/10, Writing, and NatInt. Political lean indicates ideological positioning.
Metric Base Model Gabliterated Model
Parameters (#P) 3B 3B
UGI Score 21.16 26.05
W/10 Score 4.5 6.0
Writing 9.76 13.48
NatInt 13.72 16.82
Political Lean (%) -16.9 -10.2
The Nanbeige4–3B-Thinking–2511-gabliterated model is available on Hugging Face [9]. These results across different model families and scales reinforce the robustness and general applicability of the Gabliteration technique.
A.2 Performance Preservation Bounds